OpenStudy (anonymous):

A road that is 16 miles long is represented on a map that shows a scale of 1 centimeter being equivalent to 40 kilometers. How many centimeters long does the road appear on the map? Round your result to the nearest tenth of a centimeter.

OpenStudy (anonymous):

Divide 16 miles converted to km by 40. (25.7495)/40 = 0.644 cm

OpenStudy (anonymous):

How would I solve this? Your chemistry professor mixes the contents of two beakers containing 2.3 L and 300 ml of a liquid. What is the combined amount in liters?

OpenStudy (anonymous):

2.3 + 0.3 = 2.6 liters You have to convert 300 ml to liters. 300 ml = 300/1000 liters .

OpenStudy (anonymous):

In other words, you cannot add apples and oranges.